# Quadratic Motion Polynomials With Irregular Factorizations

### v1

@@ 3 Motion Polynomials with Irregular Factorizations
As of today, not many examples of irregularly factorizable motion polynomials are known. Our aim is to find a complete description of motion polynomials of degree two with irregular factorizations and provide new examples. For this we need to answer the question how irregular factorizability can be stated algebraically. We will only be regarding monic polynomials as, by assumption, the leading coefficient is invertible and can therefore be factored out.
3.1 Conditions for Infinite Factorizability
@@
To prove the theorem we try to factorize the polynomial. Since its norm-polynomial is given by $C\tilde{C}=(t-h_{1})\widetilde{(t-h_{1})}(t-h_{2})\widetilde{(t-h_{2})}$ we can take $M\coloneqq(t-h_{2})\widetilde{(t-h_{2})}$ as a monic quadratic factor. If we now divide $C$ by $M$ we see
@@
Taking $M\coloneqq(t-h_{1})\widetilde{(t-h_{1})}$ will result in the same criterion. ∎
@@
3.2 Non-Invertibility-Condition
To complement Theorem 3.1 we will now derive an algebraic formulation of when an algebra element is non-invertible. While this is known in the community, we strive for a simplified algebraic criterion to make ensuing calculations more manageable.
@@ 3.2 Non-Invertibility-Condition
If $a$ is invertible, then $\tilde{a}$ is invertible, $(\tilde{a})^{-1}=\widetilde{(a^{-1})}$ and the inverse of $a\tilde{a}$ is $(\tilde{a})^{-1}a^{-1}$. Conversely, if $a\tilde{a}$ is invertible, then the inverse of $a$ is given by $\tilde{a}(a\tilde{a})^{-1}$. ∎
Proposition 3.3 allows us to reduce the dimensions in which we have to look for an inverse from 32 to twelve: Since $a\tilde{a}$ is its own reverse it consists only of grade 00, $1$, $4$ and $5$ blades.
@@
consisting only of grades 00, $1$, $4$ and $5$ as
@@ 3.3 Irregular Factorizations
    We prescribe a linear right motion polynomial factor $t-h_{2}$ and define $M\coloneqq(t-h_{2})\widetilde{(t-h_{2})}$.
@@
    and $t-h_{1}$ is a motion polynomial. According to [9] the latter is the case if and only if
@@
There are three conformally non-equivalent types of motions described by $t-h_{2}$ [13, 9], conformal rotation, transversion, and conformal scaling. They are distinguished by the number of real roots of the quadrance polynomial $(t-h_{2})(t-\tilde{h}_{2})$. In general we expect there to exist a quadratic motion polynomial $C=(t-h_{1})(t-h_{2})$ with irregular factorization for each pairing of these three motion types. When demonstrating this by example, we do not have to take care of the order of factors as taking the reversion of the polynomial preserves the type of motions involved but switches the motion type of the first and second factor.
@@
All solutions of the irregular factorizability condition lie on the real variety generated by the determinant of the norm of an element. Interestingly all real solutions seem so be singular points of this variety. This has been verified using the Mathematica Resolve command and checking if there exists a real non-singular point on the variety.
Indeed, there exists at least one example for each case:
@@
Seen in Figure 3(a) $h_{1}=-e_{12}+e_{13}+e_{1-}+e_{23}+e_{2-}$, $h_{2}=e_{12}$ This motion has an infinite amount of factorizations.
@@
Seen in Figure 3(b) $h_{1}=-\frac{1}{2}e_{12}+\frac{4}{5}e_{13}+\frac{49}{30}e_{1-}+\frac{4}{3}e_{1+}$, $h_{2}=e_{12}$ This motion has one unique factorization.
@@
Seen in Figure 3(c) $h_{1}=e_{13}+\frac{\sqrt{6}}{3}(2e_{1-}+1)$, $h_{2}=e_{12}$ This motion has one unique factorization.
@@
Seen in Figure 3(d) $h_{1}=e_{3-}-e_{+-}-e_{13}+e_{1+}+e_{1-}-e_{23}+e_{2+}+e_{2-}+\sqrt{2}$, $h_{2}=e_{3+}+e_{3-}$ This motion has two distinct factorizations.
@@
Seen in Figure 3(e) $h_{1}=-e_{3-}+e_{2-}+\sqrt{3}$, $h_{2}=-e_{+-}$ This motion has five distinct factorizations.
@@
Seen in Figure 3(f) $h_{1}=e_{2-}+\frac{1}{2}(\sqrt{5}e_{2+}+e_{+-})$, $h_{2}=-e_{+-}$ This motion has three distinct factorizations.
@@
Figure 3: Trajectories of points on a unit circle centered at $(0,0,3)$ under irregularly factorizable motion polynomials.
@@ 4 Rigid Body Motions
A natural question arising now is, do there exist irregularly factorizable motions in the space $\operatorname{SE}(3)$ of rigid body motions? Currently the only known non-trivial example of degree two is the circular translation given by $C\coloneqq(t-(xe_{1}+ye_{2})(e_{-}+e_{+}))(t-e_{12})$ [32]. The two factors describe rotations around parallel axes of equal speed but in opposite direction so that the resulting motion is translational. In the following we will show that this is indeed the only one.
Let us first define the notion of trivial factorizations:
@@
The motion $C=(t-h_{1})(t-h_{2})$ is called trivial if there is a map $f\colon\mathbb{R}\to\mathbb{R}$ and $h\in\mathrm{CGA}_{+}$ such that $h\tilde{h}\in\mathbb{R}$ and $C(t)=f(t)-h$.
A typical example of a trivial motion is $C=(t-h)(at+b-h)$ with $C\tilde{C}\in\mathbb{R}[t]\setminus\{0\}$ and some $a,b\in\mathbb{R}$.
Remark 4.2.
This definition essentially says that we call a motion trivial if it is a possibily non-linear reparametrization of a simple motion.
@@
Up to Euclidean equivalence the circular translation $C\coloneqq(t-(xe_{1}+ye_{2})(e_{-}+e_{+}))(t-e_{12})$ is the only non-trivial quadratic Euclidean motion that has irregular factorizations.
@@
To study Euclidean motions as special conformal motions, we reduce $\mathrm{CGA}$ to the algebra $\mathbb{D}\mathbb{H}$ of dual quaternions, given as
@@
For $(t-x)\in\mathbb{D}\mathbb{H}[t]$ to be a motion polynomial we need to enforce that $x$ lies on the Study variety and $x+\tilde{x}$ is real. Assuming $p_{i}\neq 0$, this gives us the solution $x=\frac{q_{j}p_{j}+q_{k}p_{k}}{p_{i}}e_{23}+q_{j}e_{13}-q_{k}e_{12}+p_{i}e_{1m}+p_{i}e_{1p}+p_{k}e_{3m}+p_{k}e_{3p}+p_{s}e_{123m}$. In the next step we need to specify our right factor and solve for the irregular factorizability condition (2).
@@
In conclusion there exists only trivially irregularly factorizable motions with translations as factors and only the circular translation when there is a rotation as a factor. ∎
@@ 5 Polynomials With Commuting Factors
After having investigated the special case of rigid body motions, we will now turn to the case of commuting factors. Currently, there is also only one known example which has been previously described, in [11, 12]. In this case we get the extra condition that $h_{1}h_{2}=h_{2}h_{1}$. After splitting the set of potential right factors $(t-h_{2})$ into the parts where $h_{2}$ is a rotation, translation or scaling, we look for solutions. Doing this we find the following holds true.
@@
We proceed in the same manner as in the proof of Theorem 4.3. Let $x=x_{1}+x_{2}e_{12}+x_{3}e_{13}+\cdots+x_{16}e_{23+-}$. This time we have the extra condition that $xh_{2}-h_{2}x=0$. Let us first assume $h_{2}$ to describe a rotation. Without loss of generality $h_{2}=e_{12}$. We now get two possible solutions for $x$ such that $(t-x)$ is a motion polynomial that commutes with $(t-h_{2})$.
@@
in the $n$ direction by $-d$ and observe the following properties:
@@
$ax\tilde{a}=\alpha a(p\wedge b_{2})\tilde{a}=\alpha a(p\wedge b_{2})\tilde{a}=\alpha(ap\tilde{a})\wedge(ab_{2}\tilde{a})=e_{3}\wedge(ab_{2}\tilde{a})$
@@
Let us now investigate the case of a transversion being the right factor. Without loss of generality we choose $h_{2}\coloneqq e_{3+}+e_{3-}$. Similar to the rotational case we can find two possible commuting factors $(t-x)$ and $(t-x^{\prime})$ of a motion polynomial. The infinite factorizability conditions read
@@
In the first case we get $x=1$. This generates a trivially factorizable motion. For the second case we get $x^{\prime}=x^{\prime}_{1}+(1\pm x^{\prime}_{1})e_{+-}$, which is just an offset of the original scaling. Hence also a trivial factorization. ∎

### v2

@@ 3 Motion Polynomials with Irregular Factorizations
As of today, only few examples of irregularly factorizable motion polynomials are known, the most famous of which are the circular translation [31] that is, a motion without rotational component whose trajectories are circles and the Villarceau motion [11] whose trajectories are related to Villarceau circles on a torus. Our aim is to find a complete description of motion polynomials of degree two with irregular factorizations and provide new examples. For this we need to answer the question how irregular factorizability can be stated algebraically. We will only be regarding monic polynomials as, by assumption, the leading coefficient is invertible and can therefore be factored out.
3.1 Conditions for Irregular Factorizability
@@
To prove the theorem we try to factorize the polynomial. Since its norm-polynomial is given by $C\tilde{C}=(t-h_{1})(t-h_{1}){\,\tilde{}}\;(t-h_{2})(t-h_{2}){\,\tilde{}}$ we can take $M\coloneqq(t-h_{2})(t-h_{2}){\,\tilde{}}$ as a monic quadratic factor. If we now divide $C$ by $M$ we see
@@
Taking $M\coloneqq(t-h_{1})(t-h_{1}){\,\tilde{}}$ will result in the same criterion. ∎
@@
3.2 Non-Invertibility Condition
To complement Theorem 3.1 we will now derive an algebraic formulation of when an algebra element is non-invertible. While this is known in the Geometric Algebra community, we strive for a simplified algebraic criterion to make ensuing calculations more manageable.
@@ 3.2 Non-Invertibility Condition
If $a$ is invertible, then $\tilde{a}$ is invertible, $(\tilde{a})^{-1}=(a^{-1}){\,\tilde{}}$ and the inverse of $a\tilde{a}$ is $(\tilde{a})^{-1}a^{-1}$. Conversely, if $a\tilde{a}$ is invertible, then the inverse of $a$ is given by $\tilde{a}(a\tilde{a})^{-1}$. ∎
Proposition 3.3 allows us to reduce the dimensions in which we have to look for an inverse from 32 to twelve: Since $a\tilde{a}$ is its own reverse it consists only of grade 0, 11, 44 and $5$ multivectors, which are all blades and are not changed by reversion.
@@
consisting only of grades 0, 11, 44 and $5$ as
@@ 3.3 Irregular Factorizations
    We prescribe a linear right motion polynomial factor $t-h_{2}$ and define $M\coloneqq(t-h_{2})(t-h_{2}){\,\tilde{}}$.
@@
    and $t-h_{1}$ is a motion polynomial. According to [9], the latter is the case if and only if
@@
There are three conformally non-equivalent types of motions described by $t-h_{2}$ [13, 9], conformal rotation, transversion, and conformal scaling. They are distinguished by the number of real roots of the quadrance polynomial $(t-h_{2})(t-\tilde{h}_{2})$. In general, we expect there to exist a quadratic motion polynomial $C=(t-h_{1})(t-h_{2})$ with irregular factorization for each pairing of these three motion types. When demonstrating this by example, we do not have to take care of the order of factors as taking the reversion of the polynomial preserves the type of motions involved but switches the motion type of the first and second factor.
@@
All solutions of the irregular factorizability condition lie on the real variety generated by the determinant of the quadrance of an element. Interestingly, all real solutions seem so be singular points of this variety. This has been verified using the Mathematica Resolve command and checking if there exists a real non-singular point on the variety.
Example 3.6.
There exists at least one example for each case:
@@
Seen in Figure 4(a).
$h_{1}=-e_{12}+e_{13}+e_{1-}+e_{23}+e_{2-}$,
$h_{2}=e_{12}$
This motion has an infinite amount of factorizations.
@@
Seen in Figure 4(b).
$h_{1}=-\frac{1}{2}e_{12}+\frac{4}{5}e_{13}+\frac{49}{30}e_{1-}+\frac{4}{3}e_{1+}$,
$h_{2}=e_{12}$
This motion has one unique factorization.
@@
Seen in Figure 4(c).
$h_{1}=e_{13}+\frac{\sqrt{6}}{3}(2e_{1-}+1)$,
$h_{2}=e_{12}$
This motion has one unique factorization.
@@
Seen in Figure 4(d).
$h_{1}=e_{3-}-e_{+-}-e_{13}+e_{1+}+e_{1-}-e_{23}+e_{2+}+e_{2-}+\sqrt{2}$,
$h_{2}=e_{3+}+e_{3-}$
This motion has two distinct factorizations.
@@
Seen in Figure 4(e).
$h_{1}=-e_{3-}+e_{2-}+\sqrt{3}$,
$h_{2}=-e_{+-}$
This motion has five distinct factorizations.
@@
Seen in Figure 4(f).
$h_{1}=e_{2-}+\frac{1}{2}(\sqrt{5}e_{2+}+e_{+-})$,
$h_{2}=-e_{+-}$
This motion has three distinct factorizations.
@@
Figure 4: Trajectories of points of a circle under irregularly factorizable motion polynomials.
Remark 3.7.
The graphics in Figure 4 were plotted using Asymptote. An interactive version of the figures can be reproduced using the python package cga-py [33] and plotting the trajectories for a unit circle in the $x$-$y$-Plane centered at $(0,0,3)$.
@@ 4 Rigid Body Motions
A natural question arising now is, do there exist irregularly factorizable motions in the space $\operatorname{SE}(3)$ of rigid body motions? The only known non-trivial example of degree two is the circular translation given by $C\coloneqq(t-(xe_{1}+ye_{2})(e_{-}+e_{+}))(t-e_{12})$ [34]. The two factors describe rotations around coupled parallel axes of equal speed but in opposite direction so that the resulting motion is translational, as seen in Figure 1. In the following we will show that this is indeed the only one.
Let us first define the notion of trivial factorizations. For this we want to call a motion trivial if it is a possibly non-linear reparametrization of a simple motion.
@@
The motion $C=(t-h_{1})(t-h_{2})$ is called trivial if there is a map $f\colon\mathbb{R}\to\mathbb{R}$ and $h\in\mathrm{CGA}_{+}$ such that $h\tilde{h}\in\mathbb{R}$ and C(t)=f(t)hC(t)=f(t)-h.
A typical example of a trivial motion is C=(th)(at+bh)C=(t-h)(at+b-h) with $C\tilde{C}\in\mathbb{R}[t]\setminus\{0\}$ and some $a,b\in\mathbb{R}$.
@@
Up to Euclidean equivalence the circular translation $C\coloneqq(t-((xe_{1}+ye_{2})(e_{-}+e_{+})-e_{12}))(t-e_{12})$ is the only non-trivial quadratic Euclidean motion that has irregular factorizations.
@@
To study Euclidean motions as special conformal motions, we reduce $\mathrm{CGA}$ to the algebra $\mathbb{DH}$ of dual quaternions, given as
@@
For $(t-x)\in\mathbb{DH}[t]$ to be a motion polynomial we need to enforce that $x$ lies on the Study variety and $x+\tilde{x}$ is real. Assuming $p_{i}\neq 0$, this gives us the solution $x=\frac{q_{j}p_{j}+q_{k}p_{k}}{p_{i}}e_{23}+q_{j}e_{13}-q_{k}e_{12}+p_{i}e_{1m}+p_{i}e_{1p}+p_{k}e_{3m}+p_{k}e_{3p}+p_{s}e_{123m}$. In the next step we need to specify our right factor and solve for the irregular factorizability condition (2).
@@
In conclusion, there exist only trivially irregularly factorizable motions with translations as factors and only the circular translation when there is a rotation as a factor. ∎
@@ 5 Polynomials With Commuting Factors
After having investigated the special case of rigid body motions, we will now turn to the case of commuting factors. There is also only one known example which has been previously described, in [11, 12]. In this case we get the extra condition that $h_{1}h_{2}=h_{2}h_{1}$. After splitting the set of potential right factors $(t-h_{2})$ into the parts where $h_{2}$ is a rotation, translation or scaling, we look for solutions. Doing this we find the following holds true.
@@
We proceed in the same manner as in the proof of Theorem 4.2. Let $x=x_{1}+x_{2}e_{12}+x_{3}e_{13}+\cdots+x_{16}e_{23+-}$. This time we have the extra condition that $xh_{2}-h_{2}x=0$. Let us first assume $h_{2}$ to describe a rotation. Without loss of generality $h_{2}=e_{12}$. We now get two possible solutions for $x$ such that (tx)(t-x) is a motion polynomial that commutes with $(t-h_{2})$.
@@
in the $n$ direction by d-d and observe the following properties:
@@
[…]
@@
Let us now investigate the case of a transversion being the right factor. Without loss of generality we choose $h_{2}\coloneqq e_{3+}+e_{3-}$. Similar to the rotational case we can find two possible commuting factors (tx)(t-x) and $(t-x^{\prime})$ of a motion polynomial. The infinite factorizability conditions read
@@
In the first case we get $x=1$. This generates a trivially factorizable motion. For the second case we get $x^{\prime}=x^{\prime}_{1}+(1\pm x^{\prime}_{1})e_{+-}$, which is just an offset of the original scaling. Hence, also a trivial factorization. ∎
6 Conclusion
In this text we have completely characterized quadratic motion polynomials with irregular factorizations. The restriction to polynomials of degree two allows for relatively simple computational approaches and includes already known quadratic motion polynomials. Extensions to higher degrees are of course of interest as are extensions to more general motion groups. We expect many similarities but also foresee some crucial differences, for example the algebraic description of the Study variety. As we have seen, irregularly factorizable polynomials can have infinitely many but also finitely many factorizations. We consider a characterization of the infinitely factorizable motion polynomials as an worthy topic of further research. For factorization of motion polynomials, we are interested in the decomposition of rational curves on those orthogonal groups comparing both with the 2-blades decomposition [13] of elements of those orthogonal groups. Finally, as suggested by a reviewer, it would be interesting to draw parallels to the work of Roelfs and De Keninck [35] on bivector decomposition in future work.
Data Availability Statement
No new data was created or analyzed in this study.
@@ Acknowledgments
Zijia Li is supported by the National Key R&D Program of China (2023YFA10 09401) and partially supported by the Strategic Priority Research Program of the Chinese Academy of Sciences 0640000 & XDB0640200 and the Guangdong Basic and Applied Basic Research Foundation under Grant 2024A151501 0506.